# Diffusion-SDPO: Safeguarded Direct Preference Optimization for Diffusion Models

### v1

@@ 3. Get per-branch residual objectives:
$\mathcal{L}^{w}=\tfrac{1}{2}\|\hat{\epsilon}^{w}_{\theta}-\epsilon\|_{2}^{2}-\tfrac{1}{2}\|\hat{\epsilon}^{w}_{\text{ref}}-\epsilon\|_{2}^{2},\quad\mathcal{L}^{l}=\tfrac{1}{2}\|\hat{\epsilon}^{l}_{\theta}-\epsilon\|_{2}^{2}-\tfrac{1}{2}\|\hat{\epsilon}^{l}_{\text{ref}}-\epsilon\|_{2}^{2}.$
@@
 5. Scale only loser gradients: $\mathcal{L}^{l}_{\text{scaled}}=\mathcal{L}^{l}_{\text{detach}}+\lambda_{\text{safe}}\big(\mathcal{L}^{l}-\mathcal{L}^{l}_{\text{detach}}\big)$.
 5. Mark: $\mathcal{L}^{l}_{\text{detach}}$ is a copy of $\mathcal{L}^{l}$ with gradients detached (no gradient flow).
@@
Algorithm 1 Diffusion-SDPO: Winner-preserving scaling for DPO-style diffusion training
4.3 Training with Diffusion-SDPO
At each iteration, we use preference pairs with a shared prompt $c$ and diffusion step $t$. Two forward passes yield the per-step losses $(\mathcal{L}^{w},\mathcal{L}^{l})$ and the corresponding output-space gradients $(g^{w},g^{l})$. We set the safeguard
$\lambda_{\text{safe}}=\frac{(1-\mu)\,\|g^{w}\|_{2}^{2}}{{g^{w}}^{\top}g^{l}},$ (17)
and clip it to $[0,1]$ for stability (if ${g^{w}}^{\top}g^{l}\leq 0$, we set $\lambda_{\text{safe}}=1$). Here $\mu\!\in\![0,1]$ is a safety slack to offset curvature and mini-batch noise in the first-order estimate (see Sec. B for more details). Any $\lambda<\lambda_{\text{safe}}$ ensures $\|\nabla_{\theta}\mathcal{L}^{w}\|_{2}^{2}-\lambda\,\nabla_{\theta}{\mathcal{L}^{w}}^{\top}\nabla_{\theta}\mathcal{L}^{l}>0$, hence $\Delta\mathcal{L}^{w}<0$ and the winner loss strictly decreases. For the logistic DPO objective, we implement this by scaling the backpropagated loser gradient with $\lambda_{\text{safe}}$. Algorithm 1 summarizes the procedure to integrate SDPO into Diffusion-DPO.
5 Experiments
5.1 Experimental Setting
Datasets and Models.
Following dspo; dmpo, we finetune Stable Diffusion 1.5 (SD1.5) and SDXL on preference pairs from Pick-a-Pic V2 (Pick V2) pap training set. For evaluation, we use the test prompts from Pick V2, HPS V2 hpsv2, and PartiPrompts partiprompts. Beyond SD1.5 and SDXL, we also conduct experiments on Ovis-U1 ovis-u1 (3.6B), a unified model for text-to-image synthesis and image editing. To enable DPO finetuning on Ovis-U1, we construct a mixed preference corpus that integrates text-to-image and editing pairs, totaling about 33K pairs.
Training Details and Baselines.
We integrate SDPO into Diffusion-DPO diffusion-dpo, DSPO dspo, and DMPO dmpo implementations and keep their official hyperparameters. All models are finetuned for 2000 steps with a global batch size of 2048. The learning rate is $1\times 10^{-8}$ for SD1.5 and $1\times 10^{-9}$ for SDXL. For the safeguard coefficient $\mu$, on SD 1.5 we set $0.9$ for Diffusion-DPO+SDPO and DMPO+SDPO, $0.2$ for DSPO+SDPO. On SDXL, $\mu$ is fixed as $0.6$ for all variants. We compare against several baselines: the original pretrained SD1.5 and SDXL, supervised finetuning (SFT), Diffusion-KTO diffusion-kto, MaPO mapo, DPOP dpop, and original Diffusion-DPO, DSPO, DMPO. For baselines we follow a strict hierarchy. If official checkpoints are publicly available, we evaluate those directly. If checkpoints are unavailable but official code exists, we run the released implementation with the authors’ recommended settings. If neither is available, we reimplement the method from the paper.
Evaluation.
We evaluate models on automatic preference metrics, including PickScore pap, HPS V2 hpsv2, LAION Aesthetic Classifier laion-aesthetics, CLIP clip and ImageReward imagereward scores. Sampling uses a guidance scale of 7.5 and 50 denoising steps. For Ovis-U1, we additionally evaluate structured text-to-image alignment on GenEval geneval and DPG-Bench dpg_bench, as well as image-editing performance on ImgEdit imgedit and GEdit-EN gedit.
@@ 5.1 Experimental Setting
Table 1: Reward score comparison on the HPS V2 with SD 1.5. Rows labeled “+ SDPO” report the performance obtained by applying our SDPO to the corresponding base method in the preceding row. : results from our implementation due to the lack of official code. Best results are in bold. Owing to space constraints, the full table is provided in the appendix (Table 6).
Method PickScore($\uparrow$) 0 HPS($\uparrow$) 0 Aesthetics($\uparrow$) 0 CLIP($\uparrow$) 0 Image Reward($\uparrow$)
@@
Model 1 Model 2 Pick HPS V2 Aesth. CLIP ImageReward Mean
@@ 5.2 Main Results
Table 1 shows that augmenting Diffusion-DPO, DSPO, and DMPO with SDPO consistently improves automatic reward metrics under SD 1.5, with DMPO+SDPO typically achieving the best overall scores. Win-rate results on SD 1.5 (Table 2) further confirm that each base method benefits from SDPO and that SDPO variants also outperform the SD 1.5 baseline, indicating stronger preference alignment without quality loss. On SDXL (Table 3), gains are more moderate yet consistent, suggesting reliable scaling to larger backbones. Finally, on the unified Ovis-U1 model (Table 4), SDPO yields clear improvements in preference metrics and editing scores, highlighting effectiveness for large-scale unified generation and editing. Fig. 2 compares SD 1.5 and aligned variants. While naive Diffusion-DPO can enlarge preference margins at the expense of fidelity, our safeguarded integrations preserve details and improve prompt adherence across diverse prompts. The visual evidence aligns with the quantitative trends, indicating that SDPO stabilizes optimization and enhances perceptual quality. See Sec. D for more results.
Table 3: Average win rate comparison (%) over the HPS V2 using SDXL.
Model 1 Model 2 Pick HPS V2 Aesth. CLIP ImageReward Mean
SDPO augmentation effect (base+SDPO vs base)
Diff.-DPO + SDPO Diff.-DPO 64.62 53.37 47.75 57.75 51.88 55.08
DSPO + SDPO DSPO 64.25 62.62 59.75 48.62 58.13 58.67
DMPO + SDPO DMPO 55.88 53.75 52.25 55.50 56.00 54.68
Versus SDXL
Diff.-DPO SDXL 48.25 63.38 52.50 48.38 59.13 54.33
Diff.-DPO + SDPO SDXL 59.25 66.75 49.12 55.00 58.00 57.63
DSPO SDXL 37.88 57.00 40.25 55.63 56.87 49.53
DSPO + SDPO SDXL 54.63 68.00 44.62 55.00 65.12 57.47
DMPO SDXL 58.63 64.88 46.00 53.62 62.00 57.03
DMPO + SDPO SDXL 60.88 68.50 47.63 58.38 68.63 60.80
Table 4: Comparison of Ovis-U1 ovis-u1 variants on preference, structured alignment, and image editing benchmarks. Higher is better ($\uparrow$). SDPO is particularly effective for preference alignment in large-scale models.
@@
Figure 2: Qualitative comparison of different methods using SD 1.5. Prompt: 1) The Little Prince and the fox in a Tim Burton style artwork. 2) A futuristic modern house on a floating rock island surrounded by waterfalls, moons, and stars on an alien planet. See Sec. D for more results.
@@ 5.3 Ablation Study
Table 5: Ablation study on winner-preserving rules. Model: SD 1.5. Prompts: HPS V2. $\ddagger$: fixed $\lambda_{\text{safe}}$ is used in SDPO.
Method PickScore ($\uparrow$) HPS V2 ($\uparrow$)
MaPO 0.2124 0.2760
DPOP 0.2144 0.2780
Diff.-DPO+SDPO 0.2158 0.2803
Diff.-DPO+SDPO 0.2174 0.2827
@@
Modular Ablation.
[…]
@@
Fig. 3 contrasts the training dynamics of Diff.-DPO, DSPO, and DMPO with or without SDPO. Without SDPO, $\mathcal{L}^{w}-\mathcal{L}^{l}$ decreases as expected, whereas $\mathcal{L}^{w}$ remains nondecreasing and drifts upward in Diff.-DPO and DMPO, indicating unstable optimization. With SDPO, $\mathcal{L}^{w}$ drops early and remains low, $\mathcal{L}^{l}$ declines smoothly without overshoot, and $\mathcal{L}^{w}-\mathcal{L}^{l}$ decreases steadily to a plateau. We observe a shared qualitative profile across the three SDPO-augmented settings: after basic rescaling, trajectories from different objectives largely overlap. $\mathcal{L}^{w}$ follows a monotone, fast-then-slow descent, $\mathcal{L}^{l}$ descends smoothly, and their gap grows in a stable manner across timesteps. This empirical regularity suggests that SDPO successfully corrects harmful update directions and magnitudes by acting on gradient geometry rather than on a particular objective form, thereby normalizing training dynamics across DPO variants, preserving the preferred branch, and stabilizing preference alignment.
@@
Fig. 4 shows that both HPS V2 and PickScore vary smoothly with $\mu$ and peak at moderate values, indicating low sensitivity. On SDXL, all SDPO-augmented objectives achieve near-optimal performance around $\mu\!\approx\!0.6$ with flat neighborhoods. On SD 1.5, Diffusion-DPO + SDPO and DMPO + SDPO peak near $\mu\!\approx\!0.9$, while DSPO + SDPO peaks near $\mu\!\approx\!0.2$. This aligns with Fig. 3, where DSPO already shows stable decreases in $\mathcal{L}^{w}$ and $\mathcal{L}^{l}$, requiring only a small safeguard. Therefore, by default, $\mu$ is set to 0.6 on SDXL for all objectives, while on SD 1.5 it is 0.9 for Diffusion-DPO/DMPO and 0.2 for DSPO.

### v2

@@ 3. Get per-branch residual objectives:
  -6pt
$\mathcal{L}^{w}=\tfrac{1}{2}\|\hat{\epsilon}^{w}_{\theta}-\epsilon\|_{2}^{2}-\tfrac{1}{2}\|\hat{\epsilon}^{w}_{\text{ref}}-\epsilon\|_{2}^{2},$
$\mathcal{L}^{l}\;=\tfrac{1}{2}\|\hat{\epsilon}^{l}_{\theta}-\epsilon\|_{2}^{2}-\tfrac{1}{2}\|\hat{\epsilon}^{l}_{\text{ref}}-\epsilon\|_{2}^{2}.$
  -18pt
@@
 5. Scale only loser gradients: $\mathcal{L}^{l}_{\text{scaled}}=\mathcal{L}^{l}_{\text{detach}}+\lambda_{\text{safe}}\big(\mathcal{L}^{l}-\mathcal{L}^{l}_{\text{detach}}\big)^{\dagger}$.
@@
$\dagger:\mathcal{L}^{l}_{\text{detach}}$ is a copy of $\mathcal{L}^{l}$ without gradient flow.
Algorithm 1 Training of Diffusion-SDPO.
4.2 Closed-Form Safeguard in Output Space
Directly evaluating the parameter-space bound in Eq. 14 is infeasible for a high-dimensional model, since it would require computing and storing the full gradients $\nabla_{\theta}\mathcal{L}^{w}$ and $\nabla_{\theta}\mathcal{L}^{l}$ just to take their dot product. However, we can derive a convenient proxy by considering gradients in the model’s output space. Modern diffusion models predict a noise or image tensor as output, and the training loss (e.g., a denoising score-matching loss [ddpm]) is defined on this output. Let $o^{w}$ and $o^{l}$ denote the model’s output activations for the winner and loser branches respectively (for example, $o$ could be the predicted noise residual at a certain diffusion step). Using the chain rule, we have $\nabla_{\theta}\mathcal{L}^{w}={J^{w}}^{\top}\nabla_{o}\mathcal{L}^{w}$ and $\nabla_{\theta}\mathcal{L}^{l}={J^{l}}^{\top}\nabla_{o}\mathcal{L}^{l}$, where $J$ is the Jacobian $\partial o/\partial\theta$ and $\nabla_{o}\mathcal{L}$ is the gradient of the loss with respect to the model output. Let $g^{w}=\nabla_{o}\mathcal{L}^{w}$ and $g^{l}=\nabla_{o}\mathcal{L}^{l}$ denote the output-space gradients for the winner and the loser. Eq. 14 can then be written as:
$\lambda$ $\leq\!\frac{\|\nabla_{\theta}\mathcal{L}^{w}\|_{2}^{2}}{\nabla_{\theta}{\mathcal{L}^{w}}^{\top}\nabla_{\theta}\mathcal{L}^{l}}=\frac{{g^{w}}^{\top}\!\big({J^{w}}{J^{w}}^{\top}\big){g^{w}}}{{g^{w}}^{\top}\!\big({J^{w}}{J^{l}}^{\top}\big){g^{l}}}$ (15)
$=\!\frac{\|{g^{w}}\|_{2}^{2}}{\,{g^{w}}^{\top}{g^{l}}\,}\!\cdot\!\underbrace{\frac{\,{g^{w}}^{\top}\!\big({J^{w}}{J^{w}}^{\top}\big){g^{w}}\,}{\,\|{g^{w}}\|_{2}^{2}\,}\!\!\bigg/\!\!\frac{\,{g^{w}}^{\top}\!\big({J^{w}}{J^{l}}^{\top}\big){g^{l}}\,}{\,{g^{w}}^{\top}{g^{l}}\,}}_{\rho}\!.$ (16)
The factor $\rho$ in Eq. 16 encodes local Jacobian geometry. Estimating $\rho$ during training requires parameter–space backpropagation and adds memory usage or wall–clock time (cf. Table 4). To keep the update lightweight, we do not track $\rho$ explicitly. Instead, we absorb it into a scalar safety slack $\mu\in[0,1]$ that contracts the proxy in a controlled way:
$\lambda_{\text{safe}}=\frac{(1-\mu)\,\|g^{w}\|_{2}^{2}}{{g^{w}}^{\top}g^{l}}.$ (17)
This removes any dependence on parameter–space Jacobians and uses only the output–space gradients $g^{w}$} and $g^{l}$} that are already computed for the loss, so the additional cost is negligible. The slack $\mu$ serves as a robust guardrail under arbitrary local geometry. Larger $\mu$ yields a more conservative scaling of the loser branch; smaller $\mu$ recovers a more aggressive update. In practice, we find a fixed $\mu$ works well (see Fig. 4 for ablation on $\mu$), and for an appropriate choice of $\mu$, the output-space scheme yields $\lambda_{\text{safe}}$ trajectories that closely match those obtained from parameter-space gradients (cf. Fig. 2).
During training, we clip $\lambda_{\text{safe}}$ to $[0,1]$ for stability (if ${g^{w}}^{\top}g^{l}\leq 0$, we set $\lambda_{\text{safe}}=1$). Whenever the loser’s error vector has a positive correlation with the winner’s error vector (${g^{w}}^{\top}g^{l}>0$), $\lambda_{\text{safe}}$ provides a finite limit to how strongly we can apply the loser’s gradient without risking an increase in the winner’s loss. For the logistic DPO objective, we implement this by scaling the backpropagated loser gradient with $\lambda_{\text{safe}}$. Algorithm 1 summarizes the procedure to integrate SDPO into Diffusion-DPO. For other methods, $\lambda_{\text{safe}}$ is similarly used to scale the loser branch.
@@ 4.2 Closed-Form Safeguard in Output Space
Table 1: Reward score comparison on the HPS V2 with SD 1.5. Rows labeled “+ SDPO” report the performance obtained by applying our SDPO to the corresponding base method in the preceding row. : results from our implementation due to the lack of official code. Best results are in bold. Owing to space constraints, the full table is provided in Table 7.
Method PickScore($\uparrow$) 0 HPS($\uparrow$) 0 Aes.($\uparrow$) 0 CLIP($\uparrow$) 0 IR($\uparrow$)
@@
Table 2: Reward score comparison on the HPS V2 with SDXL. : results from our implementation due to the lack of official code. The full table is provided in Table 8.
Method PickScore($\uparrow$) 0 HPS($\uparrow$) 0 Aes.($\uparrow$) 0 CLIP($\uparrow$) 0 IR($\uparrow$)
SDXL 0.2290 0.2900 6.1271 0.3847 0.9047
SFT 0.2228 0.2883 5.9689 0.3806 0.8528
MaPO 0.2293 0.2934 6.1882 0.3840 0.9703
Diff.-DPO 0.2288 0.2927 6.1380 0.3840 1.0159
+ SDPO 0.2308 0.2938 6.1284 0.3879 1.0326
DSPO 0.2273 0.2916 6.0424 0.3894 1.0054
+ SDPO 0.2293 0.2944 6.1040 0.3889 1.0745
DMPO 0.2302 0.2921 6.1101 0.3875 1.0154
+ SDPO 0.2308 0.2933 6.1113 0.3897 1.0521
@@
Model 1 Model 2 PickScore HPS V2 Aes. CLIP ImageReward Mean
@@
5 Experiments
5.1 Experimental Setting
Datasets and Models.
Following [dspo, dmpo], we finetune Stable Diffusion 1.5 (SD 1.5) and SDXL on preference pairs from Pick-a-Pic V2 (Pick V2) [pap] training set. For evaluation, we use the test prompts from Pick V2, HPS V2 [hpsv2], and PartiPrompts [partiprompts]. Beyond SD 1.5 and SDXL, we also conduct experiments on Ovis-U1 [ovis-u1] (3.6B), a DiT [dit] model trained in a unified manner to support both text-to-image synthesis and image editing. To enable DPO finetuning on Ovis-U1, we construct a mixed preference corpus that integrates text-to-image and editing pairs, totaling about 33K pairs.
Training Details and Baselines.
We integrate SDPO into Diffusion-DPO [diffusion-dpo], DSPO [dspo], and DMPO [dmpo] implementations and keep their official hyperparameters. All models are finetuned for 2000 steps with a global batch size of 2048. The learning rate is $1\times 10^{-8}$ for SD 1.5 and $1\times 10^{-9}$ for SDXL. For the safeguard coefficient $\mu$, on SD 1.5 we set $0.9$ for Diffusion-DPO+SDPO and DMPO+SDPO, $0.2$ for DSPO+SDPO. On SDXL, $\mu$ is fixed as $0.6$ for all variants. We compare against several baselines: the original pretrained SD 1.5 and SDXL, supervised finetuning (SFT), Diffusion-KTO [diffusion-kto], MaPO [mapo], DPOP [dpop], and original Diffusion-DPO, DSPO, DMPO. For baselines we follow a strict hierarchy. If official checkpoints are publicly available, we evaluate those directly. If checkpoints are unavailable but official code exists, we run the released implementation with the authors’ recommended settings. If neither is available, we reimplement the method from the paper.
Evaluation.
We evaluate models on automatic preference metrics, including PickScore [pap], HPS V2 [hpsv2], LAION Aesthetic Classifier [laion-aesthetics], CLIP [clip] and ImageReward [imagereward] scores. Sampling uses a guidance scale of 7.5 and 50 denoising steps. For Ovis-U1, we additionally evaluate structured text-to-image alignment on GenEval [geneval] and DPG-Bench [dpg_bench], as well as image-editing performance on ImgEdit [imgedit] and GEdit-EN [gedit].
@@ 5.2 Main Results
Table 1,2 show that adding SDPO to Diffusion-DPO, DSPO, and DMPO consistently improves automatic reward metrics under SD 1.5 and SDXL, with DMPO+SDPO typically giving the best overall scores. Win-rate results on SD 1.5 (Table 3) further confirm that each base method benefits from SDPO and that SDPO variants also outperform the SD 1.5 baseline, indicating stronger preference alignment without loss of quality. On SDXL, the gains are moderate yet consistent (see Table 9), suggesting reliable scaling to larger UNet [unet] backbones.
On the unified Ovis-U1 model (Table 5), SDPO yields clear improvements in preference metrics and editing scores, demonstrating effectiveness on a DiT backbone as well. While naive Diffusion-DPO can enlarge preference margins at the expense of fidelity, our safeguarded integrations preserve details and improve prompt adherence across diverse prompts. The visual evidence (cf. Fig. 578) aligns with the quantitative trends, indicating that SDPO stabilizes optimization and enhances perceptual quality.
Figure 2: Training dynamics of $\lambda_{\text{safe}}$ on SD 1.5 (left) and Ovis-U1 (right) with two computation schemes (using output-space gradients vs. parameter-space gradients). The trajectories closely match throughout training, and the output-space variant requires substantially less computation while maintaining comparable aesthetic rewards (see Table 4).
Table 4: Ablation results for winner-preserving rules on SD 1.5 (prompts: HPS V2). We report PickScore/HPS V2, one-step training Time (s) and peak GPU memory (GB) on a single NVIDIA A100 with batch size 16 and 128 gradient accumulation steps in BF16. : fixed $\lambda_{\text{safe}}$ in SDPO. : $\lambda_{\text{safe}}$ computed with parameter-space gradients.
Method PickS. ($\uparrow$) HPS ($\uparrow$) Time ($\downarrow$) GPU Mem. ($\downarrow$)
MaPO 0.2124 0.2760 162 53.8
DPOP 0.2144 0.2780 187 58.5
Diff.-DPO 0.2131 0.2743 183 57.0
Diff.-DPO+SDPO 0.2158 0.2803 183 57.0
Diff.-DPO+SDPO 0.2176 0.2828 314 63.3
Diff.-DPO+SDPO 0.2174 0.2827 184 57.1
Table 5: Comparison of Ovis-U1 [ovis-u1] variants on preference, structured alignment, and image editing benchmarks. Higher is better ($\uparrow$). SDPO is particularly effective for preference alignment in large-scale models.
@@ 5.3 Ablation Study
Figure 4: Sensitivity of SDPO to hyperparameter $\mu$ measured by HPS V2 and PickScore across SD 1.5 and SDXL on HPS V2 prompt set.
Figure 5: Qualitative comparison of different methods using SD 1.5. Prompt: 1) The Little Prince and the fox in a Tim Burton style artwork. 2) A futuristic modern house on a floating rock island surrounded by waterfalls, moons, and stars on an alien planet. See Fig. 7 for more results.
Figure 6: Comparison of Diffusion–DPO with and without SDPO under longer training on SD 1.5 (Test prompts: Pick V2).
Modular Ablation & Computational Cost.
Table 4 compares winner preserving strategies when embedded into MaPO, DPOP, and Diffusion-DPO. MaPO applies a fixed winner weight and removes the reference model, which weakens calibration of absolute error though it requires less computation and GPU memory. DPOP protects the winner through thresholded update filtering, but this rule was designed for autoregressive language models and does not fully match diffusion training dynamics. Our SDPO preserves the winner by rescaling the loser update with a safeguard coefficient $\lambda_{\text{safe}}$ selected in the output space according to directional alignment. A fixed $\lambda_{\text{safe}}$—that is, holding $\lambda_{\text{safe}}$ constant throughout training (see Table 4, row 4)—already improves over MaPO and DPOP on PickScore and HPS, whereas allowing $\lambda_{\text{safe}}$ to adapt during training yields further gains. These improvements support the hypothesis that output–space selection of $\lambda_{\text{safe}}$ stabilizes the winner while maintaining pressure to enlarge the preference margin.
Fig. 2 compares two mechanisms for computing the dynamic $\lambda_{\text{safe}}$: using output-space gradients and using parameter-space gradients. The output-space trajectory closely matches the parameter-space trajectory in both level and trend, indicating similar effectiveness. Because it reuses signals from the standard backward pass and only performs lightweight reductions, the output-space variant adds essentially no runtime or memory overhead relative to the Diffusion-DPO baseline. In contrast, as shown in Table 4, computing $\lambda_{\text{safe}}$ from parameter-space gradients requires extra vector-Jacobian products and temporary buffers for per-parameter inner products, increasing training time by about 72% and peak memory by about 11%.
Note that the two curves in Fig. 2 use different values of $\mu$, which indicates that by adjusting $\mu$, the output-space scheme can approximate the behavior of parameter-space estimation. Although computing $\lambda_{\text{safe}}$ in the output space is an approximation, our empirical results show that proper tuning of $\mu$ effectively compensates for this mismatch, allowing the scaled gradients to closely match the parameter-space formulation while retaining the same training efficiency.
@@
Fig. 3 contrasts the training dynamics of Diff.-DPO, DSPO, and DMPO with or without SDPO. Without SDPO, $\mathcal{L}^{w}-\mathcal{L}^{l}$ decreases as expected, whereas $\mathcal{L}^{w}$ remains nondecreasing and drifts upward in Diff.-DPO and DMPO, indicating unstable optimization. With SDPO, $\mathcal{L}^{w}$ drops early and remains low, $\mathcal{L}^{l}$ declines smoothly without overshoot, and $\mathcal{L}^{w}-\mathcal{L}^{l}$ decreases steadily to a plateau. For DSPO, which already regularizes branch imbalance via its score preference objective and progressively increases the weight on the winner branch, adding SDPO causes no degradation and typically yields slightly improved reward results (cf. Table 1).
We observe a shared qualitative profile across the three SDPO-augmented settings: after basic rescaling, trajectories from different objectives largely overlap. $\mathcal{L}^{w}$ follows a monotone, fast-then-slow descent, $\mathcal{L}^{l}$ descends smoothly, and their gap grows in a stable manner across timesteps. This empirical regularity suggests that SDPO successfully corrects harmful update directions and magnitudes by acting on gradient geometry rather than on a particular objective form, thereby normalizing training dynamics across DPO variants, preserving the preferred branch, and stabilizing preference alignment.
@@
Fig. 4 indicates that the safeguard slack $\mu$ admits broad, gently convex optima rather than sharp tuning. The resulting flat plateaus for both HPS V2 and PickScore suggest that SDPO’s behavior is governed primarily by gradient–alignment geometry in output space, rather than by the specific loss form. On SDXL, all SDPO–augmented objectives maintain near–optimal performance over a wide interval centered around $\mu\approx 0.6$, consistent with a smoother, higher–capacity landscape.
On SD 1.5, the more aggressive baselines (Diffusion-DPO and DMPO) benefit from a larger slack $\mu$, which further contracts the loser contribution whenever its direction conflicts with the winner. In contrast, DSPO already progressively assigns greater weight to the winner branch during training, so a smaller slack is sufficient. Practically, $\mu$ can be chosen by an early–phase heuristic: increase $\mu$ if the winner loss $\mathcal{L}^{w}$ drifts upward or oscillates; decrease $\mu$ in small steps if the preference margin stalls despite a stable $\mathcal{L}^{w}$. Following this rule of thumb, we adopt a single default on SDXL ($\mu=0.6$ for all objectives) and two defaults on SD 1.5 ($\mu=0.9$ for Diffusion–DPO and DMPO; $\mu=0.2$ for DSPO).
SDPO preserves gains with longer training.
We further examine whether the safeguarding mechanism remains effective under extended training (Fig. 6). The baseline exhibits a nonmonotonic trajectory: both HPS V2 and PickScore increase early, then plateau or decline as training proceeds. In contrast, augmenting Diffusion-DPO with SDPO yields continued gains that stabilize at a higher level, widening the gap over the baseline as steps increase. Qualitative examples in Fig. 1 corroborate this trend: baseline generations develop saturation and texture artifacts under prolonged training, whereas SDPO maintains crisp structure and consistent style. Overall, SDPO sustains improvements in preference metrics without sacrificing perceptual quality during longer training. This follows from its winner-preserving constraint, which limits the influence of loser-branch gradients and ensures that the winner loss does not increase. Without SDPO, the growing contribution of the loser branch can raise the winner loss and is often accompanied by degraded visual quality.